OpenStudy (anonymous):

What does x in the solution equal in the following system of equations? 2x-3y=12 2x-3y=4 a)3 b)-3 c)2 d)-2

OpenStudy (amistre64):

those are parallel lines.... soo...

OpenStudy (amistre64):

is there a typo in the problem?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

not 2x-3y=4 but 2x-y=4

OpenStudy (amistre64):

2x-3y =12 2x - y =4 then y = 2x -4 2x -3(2x-4) = 12 2x -6x +12 = 12 -4x = 0 x would be 0

OpenStudy (anonymous):

hmm i dont have it as an awnswer choice

OpenStudy (amistre64):

i know; check the equations to make sure you got all the ypos out for me

OpenStudy (amistre64):

ive tried it several ways and keep getting the same result x=0

OpenStudy (amistre64):

which tells me that most likely its in the equations you posted

OpenStudy (amistre64):

y = 2x -4 y= 2/3x -4 are only equal when; x=0

OpenStudy (anonymous):

at the top i mystyped 2x+ 3y=12